OpenStudy (msmr):

In a regular polygon, the ratio of the measure of an exterior angle to the measure of an interior angle is 2:13. How many sides does the polygon have?

OpenStudy (anonymous):

Exterior angle + interior angle = 180º 2x + 13x = 180 x = 12 2x = 24 Sum of interior angles = 180º(n-2) One interior angle = 180º(n-2)/n 180º(n-2)/n = 24 n = 30/13 This polygon shouldn't exist.

OpenStudy (msmr):

the answer in the back of the book is 15 sides :/ so idk...

OpenStudy (anonymous):

Oh, I'm sorry, I mixed up the ratio. x = 12 13x = 156 180(n-2)/n = 156 n = 15